OpenStudy (anonymous):

Twenty-five members of the eighth grade class at Park Center Middle School are going to a museum and then to lunch. Each student must pay an entrance fee to the museum and $7.25 for lunch. The total cost for the trip is $443.75. What is the entrance fee for one student?

OpenStudy (anonymous):

lunch is still too expensive

OpenStudy (anonymous):

btw its 6x

OpenStudy (anonymous):

(lunch cost) = 25*7.25 = ? (One entrance) = (443,75-(lunch cost))/25 =?

OpenStudy (radar):

Let x = entrance fee. then 25(x + 7.25)=$443.75 25x + $181.25 = $443.75 25x= $262.50 x=$10.50 Entrance Fee

OpenStudy (anonymous):

25 ( e + 7.25 ) = $443.75 e + 7.25 = $443.75 / 25 = $17.75 => e = $ 17.75 - 7.25 = $ 10.50